OpenStudy (anonymous):

Algebra 1 9th grade Solve each of the following equations for x. a. (FRACTIONS) x over 6 = 7 over 3 b. 3x+2=7x-8 c. (FRACTIONS) 6 over x = 4 over x+1 d. 6(x-4)=42

OpenStudy (anonymous):

\[\frac{x}{6}=\frac{7}{3}\iff 3x=6\times 7\] is as start

OpenStudy (anonymous):

second one, subtract 3x from both sides, add 8 to both sides, divide both side by 4

OpenStudy (anonymous):

\[\frac{6}{x}=\frac{4}{x+1}\] \[\frac{x}{6}=\frac{x+1}{4}\iff 6(x+1)=4x\]

OpenStudy (anonymous):

Thank you! for all your help!

OpenStudy (anonymous):

yw